In a .25 litre tube dissociation of 4 moles of NO takes place .If it`s degree of dissociation is 10% the value of kp for reaction 2NO === N2+O2 is

Grade:12

1 Answers

Shubhojit Das
5 years ago
2NO ---- N2 +O2
Degree of dissociation = moles dissociated divided by total moles present
This implies that 10 / 100 = 4 / total moles
This implies that total moles = 4 divided by 0.1 which is 40.
Therefore total moles of NO = 40 moles
2NO-------- N2 + O2
AT equilibrium 
36.             2 .       2
Therefore k c = 2 into 2 / 36 into 36 which is equal to 1 by 18 ssquar
We know that Kp = Kc (RT)∆ng
Here gelen g is zero therefore Kp = kc= 1/18^2